OpenStudy (anonymous):

√2y-1-3=0 a y=1 b y=4 c no solution d y=5 the square root is over the 2y-1

OpenStudy (anonymous):

y=5

OpenStudy (anonymous):

square both sides 2y-1 = 9 2y=10 y=5

OpenStudy (anonymous):

√2y-1-3=0 First add 3 to both sides

myininaya (myininaya):

hey if you want to extend that square root symbol you can do sqrt{2y-1}

OpenStudy (anonymous):

√2y-1=3 now sqrt booth sides 2y -1 = 9 Add 1 to both sides 2y =10 divide by 2 y =5